# SSF-Net: Spatial-Spectral Fusion Network with Spectral Angle Awareness for Hyperspectral Object Tracking

### v1

@@ Abstract
Hyperspectral video (HSV) offers valuable spatial, spectral, and temporal information simultaneously, making it highly suitable for handling challenges such as background clutter and visual similarity in object tracking. However, existing methods primarily focus on band regrouping and rely on RGB trackers for feature extraction, resulting in limited exploration of spectral information and difficulties in achieving complementary representations of object features. In this paper, a spatial-spectral fusion network with spectral angle awareness (SST-Net) is proposed for hyperspectral (HS) object tracking. Firstly, to address the issue of insufficient spectral feature extraction in existing networks, a spatial-spectral feature backbone ($S^{2}$) is designed. With the spatial and spectral extraction branch, a joint representation of texture and spectrum is obtained. Secondly, a spectral attention fusion module (SAFM) is presented to capture the intra- and inter-modality correlation to obtain the fused features from the HS and RGB modalities. It can incorporate the visual information into the HS spectral context to form a robust representation. Thirdly, to ensure a more accurate response of the tracker to the object position, a spectral angle awareness module (SAAM) investigates the region-level spectral similarity between the template and search images during the prediction stage. Furthermore, we develop a novel spectral angle awareness loss (SAAL) to offer guidance for the SAAM based on similar regions. Finally, to obtain the robust tracking results, a weighted prediction method is considered to combine the HS and RGB predicted motions of objects to leverage the strengths of each modality. Extensive experiments on the HOTC dataset demonstrate the effectiveness of the proposed SSF-Net, compared with state-of-the-art trackers.
@@ I Introduction
Visual object tracking has gained significant research attention in recent years due to its applications in autonomous driving [1], human-computer interaction [2], and intelligent security [3]. It aims to follow the same object in subsequent frames based on the initial location in the first frame of a video sequence. While RGB-based trackers have been developed to handle challenging scenarios with object density, occlusions, and long-duration tracking, they rely on visual features like color and texture, which are susceptible to background clutter. In contrast, Hyperspectral video (HSV) cameras offer the advantage of capturing both visual texture information and spectral reflectance properties of object materials [4]. This means that even when the object and background have similar colors, their material characteristics are different, which can allow for the use of material spectral reflectance to distinguish the object from the background. Furthermore, with the advancements in HS imaging equipment, acquiring hyperspectral (HS) images has become more convenient, which leads to the thriving development of HS object tracking in recent years.
@@
The deep learning-based methods [8, 9] are roughly divided into two categories: band regrouping-based methods and cross-modality fusion-based methods. The former focuses on utilizing the existing RGB trackers to deal with the HS images through transfer learning. Gao et al. [10] proposed a feature fusion network, CBFF-Net, to fuse the information from different band groups. Li et al. [11] reconstructed the HS images through the auto-encoder learning to capture the intrinsic between the bands. In this way, the HS images are regrouped into multiple three-channel false-color images. However, the band regrouping-based methods just utilize the RGB trackers to extract the feature of the regrouped HS images, which neglect the relationship between spatial-spectral features. There is a lack of feature extraction networks suitable for HS images that consider both spatial and spectral characteristics simultaneously. The latter focuses on combining the advantages of RGB trackers and HS trackers. Liu et al. [12] designed an RGB and an HS branch to form the robust representation of the objects jointly. The RGB images and HS images from the same scenario are input to the network to predict the subsequent motions of the objects. Zhao et al. [13] proposed a fusion network based on the Transformer encoder and the robust tracking results are obtained from RGB and HS streams through the self-attention mechanism. Despite these advancements, the impact of various modality fusion methods is not yet fully understood, and there remains a gap in the development of effective feature-level fusion methods for HS object tracking.
@@
Moreover, HS trackers often adopt the RGB tracking framework, which comprises a classification (CLS) head and a localization (LOC) head. However, this approach does not fully exploit the rich multi-spectral information available in HS data, leading to sub-optimal robustness in tracking performance. As depicted in Figure 1(a), the conventional tracking paradigm creates a shared feature space through encoder embeddings. The CLS and LOC heads then predict the class and location of objects respectively, utilizing depth-wise correlation computations such as generalized Euclidean or cosine metrics—yet they do not account for the spectral correlation between HS template images (TI) and HS search images (SI). As illustrated in Figure 1(b), for a trained HS tracker, although the classification and location loss converges, the included angles $\alpha$ of the TI and positive SI and $\beta$ of the TI and negative SI are uncertain. The included angles of negative pairs may be larger than positive pairs [14]. This problem makes the tracker unable to deal with the disturbance from similar negative objects. Thus, it is necessary to constrain the included angles of both of the above pairs in the common feature space, as shown in Figure 1(c).
@@
To address the above problems, a spatial-spectral fusion network (SSF-Net) with spectral angle awareness is proposed. Firstly, to extract robust features and facilitate cross-modality feature fusion, a novel spatial-spectral feature backbone $S^{2}$ especially for HS images is designed, which includes several spatial-spectral convolution blocks ($S^{2}$). Besides, with the downsample layers and residual connections of $S^{2}$, the high-level semantic information is captured and the subsequent features can be fused naturally. Secondly, to incorporate the visual information from the RGB modality into the HS branch, a spectral attention fusion module (SAFM) is proposed. Compared with the other fusion-based trackers, the SAFM captures the cross-spectral context within modalities, and the HS fused features are obtained from the RGB and the HS modalities in a self-adaptive way. So, sufficient visual prompts are supplemented to the HS branch. Finally, to fully exploit the relationships between bands and enhance the robustness of the tracking results, a spectral angle awareness module (SAAM) and a novel spectral angle awareness loss (SAAL) are proposed. Inspired by the spectral angle [15], the spectral similarity is determined by the angle between two feature vectors, which we believe can assist the classification branch in predicting the location. Between TI and SI, the spectral angle for the regions corresponding to positive samples should be small, whereas, for negative samples, it should be large. With the training of the SAAL, the SAAM is capable of accurately predicting the position of the object, serving as a complement to the CLS head. Moreover, to further improve performance, a weighted prediction module is utilized to integrate predictions from both RGB and HS modalities.
@@
    A spatial-spectral fusion network with spectral angle awareness (SST-Net) for HS object tracking is proposed, which integrates the information of the HS and RGB modalities to form the robust fusion feature representation, effectively enhancing the tracking performance.
@@
    An $S^{2}$ is proposed to extract HS features, instead of using the off-the-shelf RGB tracker, which aims to model the spatial and spectral context information jointly. Moreover, an SAFM is designed to combine the HS and RGB features with the intra- and cross-modality correlation modeling, thereby generating a more robust HS feature representation.
@@
    An SAAM is proposed to capture the spectral angle similarity between regions of feature maps to perceive the positions of objects precisely. With the guidance of SAAL, the SAAM can focus on the similar regions between TI and SI. Extensive experiments on the public dataset show the effectiveness of our proposed methods.
The remaining structure of the paper is organized as follows. Section II reviews the related works including RGB tracker, HS tracker, and fusion tracking. Section III introduces the proposed SSF-Net. Extensive ablation and comparative experimental results are shown in Section IV to show the effectiveness of the proposed methods on the HOTC datasets. Finally, Section V concludes the article.
@@ II-A Visual Object Tracker
The existing research on RGB trackers is roughly divided into two categories: trackers based on Siamese networks and trackers based on Transformers. On one hand, the Siamese network was originally proposed by SiamFC [16], which extracts features from query images and template images for similarity matching. After that, many works [17, 18] focus on extracting discriminative features and improving the robustness of tracking performance. Li et al. [19] incorporated the region proposal network into the Siamese tracker and the tracking problem is considered as a one-shot detection task. Guo et al. [20] proposed a graph attention sub-network to propagate the object information from TI to SI. With the graph attention mechanism, the object structure and part-level information can be considered. Chen et al. [21] proposed a general framework SiamBAN to simplify the complex anchor calculations. It is a unified fully convolutional network that directly predicts the category and location of objects, avoiding the anchor configurations. On the other hand, due to the excellent ability in modeling the correlation of self-attention mechanism, the Transformer-based trackers [22, 23] have attracted more attention in recent years. Ye et al. [24] designed a framework named OSTrack, which can extract the features and model relations of TI and SI in a unified way. Yan et al. proposed a Transformer-based tracker, STARK, which considers the global spatial and temporal features simultaneously. Zhu et al. [25] proposed a multi-modality tracking paradigm based on the visual prompts. With the fine-tuning of the foundation model, the Transformer-based tracker can deal with multiple modalities of images.
@@ II-B HS Object Tracker
HS imaging, with its enhanced spectral information, offers rich spectral features, which has proven to be a powerful modality in addressing the challenge of background clutter in visual object tracking. In some early works, hand-crafted features were adopted to represent objects. Nguyer et al. [26] introduced a framework that utilized radiative transfer theories to estimate object reflectance. Uzkent et al. [27] proposed a kernelized correlation filter (KCF) based tracker, which can track aerial vehicles using an adjustable cross-modality HS equipment. with the proposed KCF-in-multiple regions-of-interest (ROIs) approach, a reliable region prediction is obtained. Xiong et al. [6] proposed the material-based tracking framework to deal with the HS tracking problem. Two feature extractors are used to capture the local spectral and spatial features. Thus, the texture and spectral information can be obtained respectively. On the other hand, recent works mainly focus on using the convolutional feature to represent the object. Li et al. [28] proposed a band regrouping method to generate three-channel false-color images to leverage the RGB tracker to deal with the HS images. Li [29] adopted the pre-trained RGB tracker to predict raw tracking results and designed an image enhancement module to improve the image quality. Chen et al. [30] proposed a spectral awareness module to regroup the bands considering the spectral nonlinear interactions and the dynamic TI matching strategy generated reliable tracking results. Islam et al. [31] proposed a background-aware band selection method that captures the spatial changes of each band. The most important bands can be selected to generate the false-color images.
@@ II-C Fusion Tracking
Multi-modality fusion tracking, which integrates information from different modalities, has become a prominent area of focus in object tracking. One modality combination that has garnered significant attention is the fusion of RGB and infrared modalities. This approach has proven to be highly effective in enhancing tracking performance, especially in challenging conditions. Zhang et al. [32] proposed a discriminative model prediction tracker (DMP), which adopted a multi-level fusion mechanism and presented an object estimation network to improve the performance of RGB-T tracking. Zhang et al. [33] proposed an RGB-T tracker JMMAC, which considered the appearance and motion cues of objects jointly to generate reliable features. The modality fusion weights are obtained from offline-trained networks, avoiding training costs. Lan et al. [34] proposed a graph-based label prediction model to capture the relationship of the samples and obtain the importance weights of modalities. With the measure of modality discrimination, the model can predict robust tracking results. As for HS object tracking, the existing work has not yet given sufficient attention to this aspect. Liu et al. [12] proposed a fusion-based tracker SiamHYPER, which extracted both RGB and HS features. With the channel attention mechanism [35], the fused features are obtained. In the inference stage, the RGB and HS branches are used to predict the motion of objects respectively. Zhao et al. [13] proposed a Transformer-based network named TFTN, which captured the cross-modality and intra-modality information of different modalities.
@@ III-A Framework Overview
The overall network structure is shown in Figure 2. Firstly, an RGB stream and an HS stream are used to extract features from the input image. To address the issue of insufficient modeling of HS images using existing methods, $S^{2}$ is proposed for feature extraction of HS streams. The $S^{2}$ includes the spatial-spectral convolutional block ($S^{2}$), which effectively captures both spatial and spectral information, resulting in more discriminative features for cross-modality fusion tracking. The RGB stream adopts the conventional Res-Net architecture and serves as a backbone network. It is worth noting that the two streams of the SSF-Net backbone network also have four stages. It can simultaneously extract features from different semantic levels, and it facilitates feature fusion with the RGB-stream backbone network at the same stage. In addition, an SAFM is proposed for spectral feature fusion of RGB-stream and HS-stream. Specifically, considering that different bands (spectra) contain different semantic information, a spectral attention network is proposed to enhance cross-modality feature representation, which flows from the backbone network to SAFM in a gating mechanism, then performs cross-modality feature fusion to obtain HS fused features. Then, to make robust tracking results, we simultaneously predict the object motions on both RGB and HS streams and adopt a weighted ensemble strategy to fuse the tracking results of the two streams. The RGB stream utilizes the tracking header of the existing object tracking framework [21], thus the classification and location maps can be obtained. For the HS stream, the SAAM is proposed, which consists of a spectral angle awareness branch, along with classification and localization branches. Inspired by spectral angle mapping [15], this module calculates the similarity between spectra by measuring the angle between spectral feature vectors. By comparing spectral angles between different regions, the tracker can identify similar regions between two frames. Finally, to effectively train the network, the SAAL is proposed to increase the predicted spectral similarity between the TI and object regions of SI and lead the tracker to focus on the similar regions between TI and SI. After that, the SAAL is combined with classification loss and location loss for multi-task training of the network, which ensures that the network learns to effectively utilize spectral information for accurate tracking.

### v2

@@ Abstract
Hyperspectral video (HSV) offers valuable spatial, spectral, and temporal information simultaneously, making it highly suitable for handling challenges such as background clutter and visual similarity in object tracking. However, existing methods primarily focus on band regrouping and rely on RGB trackers for feature extraction, resulting in limited exploration of spectral information and difficulties in achieving complementary representations of object features. In this paper, a spatial-spectral fusion network with spectral angle awareness (SSF-Net) is proposed for hyperspectral (HS) object tracking. Firstly, to address the issue of insufficient spectral feature extraction in existing networks, a spatial-spectral feature backbone ($S^{2}$) is designed. With the spatial and spectral extraction branch, a joint representation of texture and spectrum is obtained. Secondly, a spectral attention fusion module (SAFM) is presented to capture the intra- and inter-modality correlation to obtain the fused features from the HS and RGB modalities. It can incorporate the visual information into the HS context to form a robust representation. Thirdly, to ensure a more accurate response to the object position, a spectral angle awareness module (SAAM) is designed to investigate the region-level spectral similarity between the template and search images during the prediction stage. Furthermore, a novel spectral angle awareness loss (SAAL) is developed to offer guidance for the SAAM based on similar regions. Finally, to obtain the robust tracking results, a weighted prediction method is considered to combine the HS and RGB predicted motions of objects to leverage the strengths of each modality. Extensive experiments on the HOTC-2020, HOTC-2024, and BihoT datasets demonstrate the effectiveness of the proposed SSF-Net compared with state-of-the-art trackers. The source code will be available at https://github.com/hzwyhc/hsvt.
@@ I Introduction
Visual object tracking has gained significant research attention in recent years due to its applications in autonomous driving [1, 2], human-computer interaction [3], and intelligent security [4, 5]. It aims to follow the same object in subsequent frames based on its initial location in the first frame of a video sequence. The traditional RGB tracking methods exhibit poor robustness in complex scenarios, such as lighting variations, occlusion, and background interference, making them prone to object drift or tracking failure. Moreover, traditional RGB tracking struggles to effectively distinguish objects from the background when their visual features are similar, such as in camouflaged objects or low-contrast scenes. In contrast, hyperspectral object tracking (HOT) leverages hyperspectral (HS) imaging technology, which captures object information across multiple continuous spectral bands (typically ranging from dozens to hundreds) spanning from visible light to infrared. This enables each pixel to possess rich spectral features, allowing HS data to reveal the physical and chemical composition of objects [6]. Consequently, HOT can effectively track objects even in shadows, occlusions, or weak-textured backgrounds. Furthermore, it can utilize spectral-spatial joint modeling, enhancing its adaptability to complex background interference. Therefore, compared to traditional RGB tracking, HOT offers stronger object discrimination and better environmental adaptability, making it a more robust and effective solution in challenging tracking scenarios.
@@
The deep learning-based methods [10, 11] are roughly divided into two categories: band regrouping-based methods and cross-modality fusion-based methods. The former focuses on utilizing the existing RGB trackers to deal with the HS images through transfer learning. Gao et al. [12] proposed a feature fusion network, CBFF-Net, to fuse the information from different band groups. Li et al. [13] adopted the auto-encoder learning to capture the intrinsic spectral information for the HS images to be grouped into multiple three-channel false-color images. However, the band regrouping-based methods just utilize the RGB trackers to extract features from the regrouped HS images, which neglects the relationship between spatial-spectral features. There is a lack of feature extraction networks suitable for HS images that consider both spatial and spectral characteristics simultaneously. The latter focuses on combining the advantages of RGB trackers and HS trackers. Liu et al. [14] designed an RGB and an HS branch to form the robust representation of objects jointly. The RGB and HS images from the same scenario are input to the network to predict the subsequent motions. Zhao et al. [15] proposed a fusion network based on the Transformer encoder, and the robust tracking results are obtained from RGB and HS streams through the self-attention mechanism. Despite these advancements, the impact of various modality fusion methods is not yet fully understood, and there remains a gap in the development of effective feature-level fusion methods for HS object tracking.
@@
Moreover, HS trackers often adopt an RGB tracking framework, which comprises a classification (CLS) head and a localization (LOC) head. However, such an approach does not fully exploit the rich spectral information in HS data, leading to sub-optimal tracking performance. As depicted in Figure 1(a), the conventional tracking paradigm creates a shared feature space through encoder embeddings. The CLS and LOC heads then predict the class and location of objects, respectively, utilizing depth-wise correlation computations such as generalized Euclidean or cosine metrics—yet they do not account for the spectral correlation between HS template images (TI) and HS search images (SI). As illustrated in Figure 1(b), for a trained HS tracker, although the classification and location loss converge, the included angles $\alpha$ of the TI and positive SI and $\beta$ of the TI and negative SI are uncertain. The included angles of negative pairs may be larger than positive pairs [16]. This problem makes the tracker unable to deal with the disturbance from similar negative objects. Thus, it is necessary to constrain the included angles of both of the above pairs in the common feature space, as shown in Figure 1(c).
@@
To address the above problems, a spatial-spectral fusion network (SSF-Net) with spectral angle awareness is proposed. Firstly, to extract robust features and facilitate cross-modality feature fusion, a novel spatial-spectral feature backbone $S^{2}$, especially for HS images, is designed, which includes several spatial-spectral convolution blocks ($S^{2}$). Besides, with the downsample layers and residual connections of $S^{2}$, the high-level semantic information is captured, and the subsequent features can be fused naturally. Secondly, to incorporate the visual information from the RGB modality into the HS branch, a spectral attention fusion module (SAFM) is proposed. Compared with other fusion-based trackers, the SAFM captures the cross-spectral context within modalities, and the HS fused features are obtained from the RGB and the HS modalities in a self-adaptive way. Consequently, sufficient visual prompts are supplemented to the HS branch. Finally, to fully exploit the relationships between bands and enhance the robustness of the tracking results, a spectral angle awareness module (SAAM) and a novel spectral angle awareness loss (SAAL) are proposed. Inspired by the spectral angle [17], the spectral similarity is determined by the angle between two feature vectors, which can assist the classification branch in location prediction. Between TI and SI, the spectral angle for the regions corresponding to positive samples should be small, whereas for negative samples, it should be large. With the training of the SAAL, the SAAM is capable of accurately predicting the object position, serving as a complement to the CLS head. Moreover, to further improve performance, a weighted prediction module is utilized to integrate predictions from both RGB and HS modalities.
@@
    A spatial-spectral fusion network with spectral angle awareness (SSF-Net) for HS object tracking is proposed, which integrates the information of the HS and RGB modalities to form the fused feature representation, effectively enhancing the tracking performance.
@@
    An $S^{2}$ is proposed to extract HS features instead of using the off-the-shelf RGB tracker, which aims to model the spatial and spectral context information jointly. Moreover, an SAFM is designed to combine the HS and RGB features with the intra- and inter-modality correlation modeling, thereby generating a more robust HS feature representation.
@@
    An SAAM is proposed to capture the spectral angle similarity between regions of feature maps to perceive the positions of objects accurately. With the guidance of SAAL, the SAAM can focus on the similar regions between TI and SI. Extensive experiments on public datasets demonstrate the effectiveness of our proposed methods.
The remaining paper is organized as follows. Section II reviews the related works, including RGB tracker, HS tracker, and fusion tracking. Section III introduces the proposed SSF-Net. Extensive ablation and comparative experimental results are shown in Section IV to show the effectiveness of the proposed methods on the HOTC-2020, HOTC-2024, and BihoT [18] datasets. Finally, Section V concludes the article.
@@ II-A Visual Object Tracker
The existing research on RGB trackers is roughly divided into two categories: trackers based on Siamese networks and trackers based on Transformers. The Siamese network was originally proposed by SiamFC [19], which extracts features from query images and template images for similarity matching. After that, many works [20, 21] focus on extracting discriminative features and improving the robustness of tracking performance. Li et al. [22] incorporated the region proposal network into the Siamese tracker, and the tracking problem is considered as a one-shot detection task. Guo et al. [23] proposed a graph attention sub-network to propagate the object information from TI to SI. With the graph attention mechanism, the object structure and part-level information can be considered. Borsuk et al. [24] proposed a fast, efficient, accurate, and robust Siamese visual tracker through a dual-template representation for object model adaptation. Kang et al. [25] designed a lightweight bridge module to incorporate the high-level information of deep features into the shallow large-resolution features, which produces better features for the tracking head. In Ref. [26], a Siamese region proposal network with a Gaussian head is proposed for object tracking. Chen et al. [27] proposed a general framework, SiamBAN, to simplify the complex anchor calculations, which is a unified fully convolutional network that directly predicts the category and location of objects, avoiding the anchor configurations. On the other hand, due to the excellent ability to model the correlation of the self-attention mechanism, the Transformer-based trackers [28, 29] have attracted more attention in recent years. Ye et al. [30] designed a framework named OSTrack, which can extract the features and model relations of TI and SI in a unified way. Blatter et al. [31] introduced a lightweight exemplar Transformer, which utilizes an attention layer based on a single instance for real-time object tracking. Yan et al. [32] proposed a Transformer-based tracker, STARK, which considers the global spatial and temporal features simultaneously. Zhu et al. [33] proposed a multi-modality tracking paradigm based on visual prompts. With the fine-tuning of the foundation model, the Transformer-based tracker can deal with multiple modalities of images.
@@ II-B HS Object Tracker
HS imaging, with its enhanced spectral information, offers rich spectral features, which have proven to be a powerful modality in addressing the challenge of background clutter in visual object tracking. In some early works, hand-crafted features were adopted to represent objects. Nguyer et al. [34] introduced a framework that utilized radiative transfer theories to estimate object reflectance. Uzkent et al. [35] proposed a kernelized correlation filter (KCF) based tracker, which can track aerial vehicles using an adjustable cross-modality HS equipment. With the proposed KCF-in-multiple regions-of-interest (ROIs) approach, a reliable region prediction is obtained. Xiong et al. [8] proposed the material-based tracking framework to deal with the HS tracking problem. Two feature extractors are used to capture the local spectral and spatial features. Thus, texture and spectral information can be obtained respectively. On the other hand, recent works mainly focus on using convolutional features to represent the object. Li et al. [36] proposed a band regrouping method to generate three-channel false-color images to leverage the RGB tracker to deal with the HS images. Li et al. [37] adopted the pre-trained RGB tracker to predict raw tracking results and designed an image enhancement module to improve the image quality. Chen et al. [38] proposed a spectral awareness module to regroup the bands considering the spectral nonlinear interactions, and the dynamic TI matching strategy generated reliable tracking results. Islam et al. [39] proposed a background-aware band selection method that captures the spatial changes of each band. The most important bands can be selected to generate the false-color images.
@@ II-C Fusion Tracking
Multi-modality fusion tracking, which integrates information from different modalities, has become a prominent area of focus in object tracking. One modality combination that has garnered significant attention is the fusion of RGB and infrared modalities. This approach has proven to be highly effective in enhancing tracking performance, especially in challenging conditions. Zhang et al. [40] proposed a discriminative model prediction tracker (DMP), which adopted a multi-level fusion mechanism and presented an object estimation network to improve the performance of RGB-T tracking. Zhang et al. [41] proposed an RGB-T tracker, JMMAC, which considered the appearance and motion cues of objects jointly to generate reliable features. The modality fusion weights are obtained from offline-trained networks, avoiding training costs. Lan et al. [42] proposed a graph-based label prediction model to capture the relationship of the samples and obtain the importance weights of modalities. With the measure of modality discrimination, the model can predict robust tracking results. As for HS object tracking, the existing work has not yet given sufficient attention to this aspect. Liu et al. [14] proposed a fusion-based tracker SiamHYPER, which extracted both RGB and HS features. With the channel attention mechanism [43], the fused features are obtained. In the inference stage, the RGB and HS branches are used to predict the motion of objects. Zhao et al. [15] proposed a Transformer-based network named TFTN, which captured the cross-modality and intra-modality information of different modalities.
@@ III-A Framework Overview
The overall network structure is shown in Figure 2. Firstly, an RGB stream and an HS stream are used to extract features from the input image. To address the issue of insufficient modeling of HS images using existing methods, $S^{2}$ is proposed for feature extraction of HS streams. The $S^{2}$ includes the spatial-spectral convolutional block ($S^{2}$), which effectively captures both spatial and spectral information, resulting in more discriminative features for cross-modality fusion tracking. The RGB stream adopts the conventional Res-Net architecture and serves as a backbone network. It is worth noting that the two streams of the SSF-Net backbone network also have four stages. It can simultaneously extract features from different semantic levels, and it facilitates feature fusion with the RGB-stream backbone network at the same stage. In addition, an SAFM is proposed for spectral feature fusion of the RGB-stream and the HS-stream. Specifically, considering that different bands (spectra) contain different semantic information, a spectral attention network is proposed to enhance cross-modality feature representation, which flows from the backbone network to SAFM in a gating mechanism and then performs cross-modality feature fusion to obtain HS fused features. Then, for robust tracking, the object motions on both RGB and HS streams are simultaneously predicted, and a weighted ensemble strategy is adopted to fuse the tracking results of the two streams. The RGB stream utilizes the tracking head of the existing object tracking framework [27], thus the classification and location maps can be obtained. For the HS stream, the SAAM is proposed, which consists of a spectral angle awareness branch, along with classification and localization branches. The mixed template features and search features from RGB and HS backbones through SAFM are then fed into SAAM for object position prediction. Inspired by spectral angle mapping [17], this module calculates the similarity between spectra by measuring the angle between spectral feature vectors. By comparing spectral angles between different regions, the tracker can identify similar regions between two frames. Finally, to effectively train the network, the SAAL is proposed to increase the predicted spectral similarity between the TI and object regions of SI and lead the tracker to focus on the similar regions between TI and SI. After that, the SAAL is combined with classification loss and location loss for multi-task training of the network, which ensures that the network learns to effectively utilize spectral information for accurate tracking.